# Context-Aware Deep Learning for Robust Channel Extrapolation in Fluid Antenna Systems

### v2

@@ Abstract
Fluid antenna systems (FAS) offer remarkable spatial flexibility but face significant challenges in acquiring high-resolution channel state information (CSI), leading to considerable overhead. To address this issue, we propose CANet, a robust deep learning model for channel extrapolation in FAS. CANet combines context-adaptive modeling with a cross-scale attention mechanism and is built on a ConvNeXt v2 backbone to improve extrapolation accuracy for unobserved antenna ports. To further enhance robustness, we introduce a novel spatial amplitude perturbation strategy, inspired by frequency-domain augmentation techniques in image processing. This motivates the incorporation of a Fourier-domain loss function, capturing frequency-domain consistency, alongside a spectral structure consistency loss that reinforces learning stability under perturbations. Our simulation results demonstrate that CANet outperforms benchmark models across a wide range of signal-to-noise ratio (SNR) levels.
@@ I Introduction
With the advent of sixth-generation (6G) mobile communications, the demand for higher capacity, wider coverage, and greater flexibility in wireless systems is accelerating. Meeting these ambitious requirements necessitates the development of advanced physical-layer technologies. Among them, the fluid antenna system (FAS) [1, 2] has emerged as a promising antenna architecture that introduces new spatial degrees of freedom (DoF) into wireless communication. FAS represents the capability of shape-and-position-reconfigurable antenna technology to empower the physical layer. By leveraging mechanical motion [3], liquid metals [4], reconfigurable pixel-based structures [5] or metamaterials [6], FAS enables dynamic port selection from a massive number of predefined locations. This adaptability facilitates spatial diversity gains in nearly continuous space and enhances the system’s capability to respond intelligently to the wireless channel. In [7], Lu et al. explained FAS from the viewpoint of electromagnetics.
Since the introduction of FAS by Wong et al. in [8, 9], there have been lots of efforts in understanding the achievable performance of FAS [10, 11]. The use of FAS has also been considered for reconfigurable intelligent surface (RIS) systems [12], index modulation [13, 14], wideband communication [15], and integrated sensing and communication (ISAC) [16]. Capacity maximization for FAS channels has also been studied using statistical channel state information (CSI) [17] and instantaneous CSI [18]. The diversity-multiplexing trade-off for multiple-input multiple-output (MIMO) FAS channels was analyzed in [19].
@@
Nonetheless, the promising performance of FAS relies on the availability of CSI and the high-resolution nature of FAS also leads to substantial overhead in acquiring CSI, including pilot overhead, feedback latency, and energy consumption[20]. In recent years, several studies explored the integration of physical modeling and data-driven approaches for CSI prediction [21, 22]. For instance, [23] proposed an inference algorithm to enable high-accuracy extrapolation in a massive FAS using only a small number of observed ports. Additionally, in [24], an extrapolation scheme based on masked language models, which reconstructs the full channel from a limited number of observed port CSI, was proposed. Going one step further, the approach in [25] integrates the diffusion mechanism of graph neural networks with an asymmetric masked autoencoder architecture, significantly improving both extrapolation accuracy and flexibility in high-resolution FAS settings. Prior to these, there were attempts that exploited channel sparsity [26] and used Bayesian approaches [27]. CSI prediction for multiuser FAS using deep learning was also reported in [28]. Overall, however, reliable CSI acquisition seems only possible at high signal-to-noise ratio (SNR) and it is important to study enhancement techniques that work well in noisy situations.
Motivated by the above, in this correspondence, we propose CANet, a novel deep learning framework designed for channel extrapolation in FAS. CANet adaptively infers the CSI of the unobserved ports with varying positions within a FAS, and maintains robust performance across different SNR scenarios. The architecture of CANet incorporates a context adaptive block (CAB), a cross-scale contextual attention (CSCA) mechanism, and a ConvNeXt v2 backbone enhanced with dropout regularization. This combination empowers CANet to process unobserved port positions both locally and globally, significantly improving the extraction of CSI features. To enhance model robustness further, we also introduce a spatial amplitude perturbation strategy. Besides, we refine the loss function by incorporating a fast Fourier transform (FFT) loss, which collaborates with a spectral structure consistency loss to reinforce model stability. Numerical results indicate that CANet surpasses benchmark models, demonstrating superior channel extrapolation robustness across varying noise levels.
@@ II Channel Extrapolation for FAS
Consider a downlink system where a base station (BS) with $M_{t}$ fixed-position antennas (FPAs) communicates with $U_{t}$ user terminals on the same time-frequency resource unit. Each user terminal employs a 2D FAS with $N_{s}=N_{x}\times N_{y}$ switchable positions (or ports), which are evenly distributed over a planar surface of physical dimensions $W_{s}=W_{x}\times W_{y}$. In a 3D rich scattering environment, the spatial correlation between port $(n_{x}^{s},n_{y}^{s})$ and port $(\tilde{n}_{x}^{s},\tilde{n}_{y}^{s})$ is given by
$\begin{array}[]{l}J_{\left(n_{x},n_{y}\right),\left(\tilde{n}_{x},\tilde{n}_{y}\right)}\\ \quad=j_{0}\left(\frac{2\pi}{\lambda}\sqrt{\left(\frac{\left|n_{x}-\tilde{n}_{x}\right|}{N_{x}-1}W_{x}\right)^{2}+\left(\frac{\left|n_{y}-\tilde{n}_{y}\right|}{N_{y}-1}W_{y}\right)^{2}}\right),\end{array}$ (1)
where $j_{0}(\cdot)$ denotes the spherical Bessel function of the first kind, which can be approximated by $\mathrm{sinc}(x)={\sin x}/{x}$, and $\lambda$ represents the carrier wavelength. The downlink channel for a typical user can then be modeled as
$\mathbf{g}=\delta\mathbf{U}_{s}\sqrt{\mathbf{\Lambda}_{s}}\mathbf{G}+\mathbf{n},$ (2)
where $\mathbf{U}_{s}\in\mathbb{C}^{N_{s}\times N_{s}}$ and $\mathbf{\Lambda}_{s}\in\mathbb{C}^{N_{s}\times N_{s}}$ are defined in accordance with [19], representing the eigenvector matrix and eigenvalue diagonal matrix obtained from the eigen-decomposition of the spatial correlation matrix $\mathbf{J}_{s}=\mathbf{U}_{s}\mathbf{\Lambda}_{s}\mathbf{U}_{s}^{H}$ , whose entries are given by (1). The matrix $\mathbf{G}\in\mathbb{C}^{N_{s}\times M_{t}}$consists of independent complex Gaussian random variables with zero mean and variance of $1/2$. The parameter $\delta$ characterizes the path loss, and ${\bf n}$ is the additive white Gaussian noise.
III Proposed Approach
@@ III-A Masking Strategy
We denote the perfect CSI as $\mathbf{g}\in\mathbb{C}^{N_{s}\times M_{t}}$, and construct a tensor $\mathbf{u}\in\mathbb{R}^{2M_{t}\times N_{y}\times N_{x}}$ to represent the concatenated real and imaginary components of the CSI across a 2D fluid antenna grid of size $N_{x}\times N_{y}$. To simulate partial observations, a subset of antenna ports is randomly masked along the spatial grid, with a masking ratio typically ranging from $80\%$ to $95\%$. The resulting masked tensor is denoted as $\tilde{\mathbf{u}}\in\mathbb{R}^{2M_{t}\times N_{y}\times N_{x}}$, in which the unobserved positions are replaced with a distinct negative value (e.g., $-10$) to make them distinguishable for the network. Additionally, we define a binary indicator matrix $\mathbf{E}_{\text{flag}}\in\{0,1\}^{N_{y}\times N_{x}}$, where a value of $1$ indicates an observed port and 00 is a masked (unobserved) port. No prior knowledge of the channel is used to select ports for extrapolation. Instead, a random masking strategy is employed to ensure flexibility and enhance the generalization capability of the model.
@@ III-B Model Architecture
The overall architecture of CANet consists of four training modules. Next, we summarize the role of each module.
@@ III-B1 CAB
For channel extrapolation tasks with strong spatial correlations, using convolutional neural network is a natural choice for feature extraction. However, conventional convolution is susceptible to interference from unobserved ports within the receptive field, which can degrade the quality of feature representations. To overcome this issue, we propose a dual-branch architecture that separately models and predicts the CSI of observed and unobserved ports, thereby enhancing both the modeling accuracy and the robustness of extrapolation.
@@
[…]
$m\in\sum_{i\in\mathcal{B}_{o}}w_{i}^{c}x_{i}.$ (5)
@@ III-B2 CSCA
In the CSCA module, we first reduce the input feature channels to half of the original before feeding them into two self-attention modules. One branch directly undergoes self-attention at the same scale, while the other duplicates the feature map and down-samples it to half of the original resolution. The downsampled feature map is then processed along with the original-scale features within the self-attention mechanism. At each self-attention stage, patches are extracted from the input features, and the cosine similarity between patches in the missing region and those in the external region is computed to establish cross-region feature correlations. Formally, let $q_{i}$ denote the $i$-th patch from the observed region $\psi^{l}$, and $k_{j}$ denote the $j$-th patch from the masked region $\phi^{l}$. The cosine similarity is computed as
@@
TABLE I: Architecture of the proposed CANet
Layer Operator Input Size Output Size Output Channel Kernel Size Stride Nonlinearity
1 CAB $32~{}\times~{}16$ $32~{}\times~{}16$ $64$ $5$ $1$ LeakyReLU
2 CAB $32~{}\times~{}16$ $32~{}\times~{}16$ $128$ $3$ $2$ LeakyReLU
3 CAB $32~{}\times~{}16$ $16~{}\times~{}8$ $128$ $3$ $1$ LeakyReLU
4 CAB $16~{}\times~{}8$ $16~{}\times~{}8$ $256$ $3$ $1$ LeakyReLU
5 CAB $16~{}\times~{}8$ $16~{}\times~{}8$ $256$ $3$ $1$ LeakyReLU
6 CSCA $16~{}\times~{}8$ $16~{}\times~{}8$ $256$ Null Null Null
7–16 ConvNeXt v2 $16~{}\times~{}8$ $16~{}\times~{}8$ $256$ $7$ $3$ GeLU
17 BilinerUp $16~{}\times~{}8$ $32~{}\times~{}16$ $256$ $3$ $1$ Null
18 CAB $32~{}\times~{}16$ $32~{}\times~{}16$ $128$ $3$ $1$ LeakyReLU
19 CAB $32~{}\times~{}16$ $32~{}\times~{}16$ $128$ $3$ $1$ LeakyReLU
20 CAB $32~{}\times~{}16$ $32~{}\times~{}16$ $64$ $3$ $1$ LeakyReLU
21 CAB $32~{}\times~{}16$ $32~{}\times~{}16$ $64$ $3$ $1$ LeakyReLU
22 Conv $32~{}\times~{}16$ $32~{}\times~{}16$ $16$ $7$ $1$ Null
@@ III-B3 ConvNeXt v2 with dropout regularization
After CAB and CSCA, we adopt ConvNeXt V2 [29] with dropout regularization as the backbone network to further enhance feature extraction and improve generalization. ConvNeXt has been widely used for its strong feature representation capabilities, and ConvNeXt V2 builds upon this by introducing the global response normalization (GRN) module, which enables more expressive feature representations. In our design, we employ the ConvNeXt V2 block for channel extrapolation. By randomly dropping a subset of neurons during training, dropout alters the effective network structure in each training epoch, leading to a degree of statistical independence across different training instances, thereby reducing prediction variance and enhancing generalization under diverse noise conditions.
@@ III-B4 Spatial amplitude perturbation
To enhance the model’s robustness, we propose a spatial amplitude perturbation strategy at the input stage, inspired by frequency perturbation techniques in the image domain [30, 31]. Specifically, the input CSI features are first zero-mean normalized to remove the DC component, then transformed into the frequency domain via FFT. The amplitude spectrum is then perturbed in a stochastic manner based on a predefined probability, effectively simulating potential frequency interference and noise. This approach improves the model’s adaptability to unseen distributions:
@@
where $\tilde{\mathbf{z}}$ is the spatially zero-mean feature, $\hat{\mathbf{z}}$ is the final perturbed feature after reconstruction, $\mathcal{F}(\cdot)$ and $\mathcal{F}^{-1}(\cdot)$ denote the 2D discrete Fourier transform and its inverse. The perturbation mask $\bm{\delta}\in\{0,1\}^{H\times W}$ above is sampled from a Bernoulli distribution with probability $\mu$, and the amplitude coefficients $\bm{\xi}\in[0,1]^{H\times W}$ are sampled from a normal distribution $\mathcal{N}(0,1)$. The hyperparameters $\mu$ and $\gamma$ control the perturbation probability and maximum strength, respectively. Only the amplitude of each frequency component is perturbed, while the phase is fixed to maintain structural consistency.

### v4

@@ Abstract
Fluid antenna systems (FAS) offer remarkable spatial flexibility but face significant challenges in acquiring high-resolution channel state information (CSI), leading to considerable overhead. To address this issue, leveraging the inherent spatial continuity of electromagnetic fields, we propose CANet, a robust deep learning model for channel extrapolation in FAS. CANet combines context-adaptive modeling with a cross-scale attention mechanism and is built on a ConvNeXt v2 backbone to improve extrapolation accuracy for unobserved antenna ports. To further enhance robustness, we introduce a novel spatial amplitude perturbation strategy, inspired by frequency-domain augmentation techniques in image processing. This motivates the incorporation of a Fourier-domain loss function, capturing frequency-domain consistency, alongside a spectral structure consistency loss that reinforces learning stability under perturbations. Our simulation results demonstrate that CANet not only outperforms benchmark models in terms of estimation accuracy across a wide range of signal-to-noise ratio (SNR) levels and carrier frequencies, but also significantly reduces the outage probability with moderate computational complexity.
@@ I Introduction
With the advent of sixth-generation (6G) mobile communications, the demand for higher capacity, wider coverage, improved positioning and greater flexibility in wireless systems is accelerating[34, 7, 6, 3]. Meeting these ambitious requirements necessitates the development of advanced physical-layer technologies. Among them, the fluid antenna system (FAS) [30, 20, 9] has emerged as a promising antenna architecture that introduces new spatial degrees of freedom (DoF) into wireless communication. FAS represents the capability of shape-and-position-reconfigurable antenna technology to empower the physical layer. By leveraging mechanical motion [2], liquid metals [25], reconfigurable pixel-based structures [38] or metamaterials [17, 19], FAS enables dynamic port selection from a massive number of predefined locations. This adaptability facilitates spatial diversity gains in nearly continuous space and enhances the system’s capability to respond intelligently to the wireless channel. Since the introduction of FAS by Wong et al. in [28, 29], there have been lots of efforts in understanding the achievable performance of FAS [22, 24], showing potential for reconfigurable intelligent surface (RIS) systems [15], index modulation [40, 41], wideband communication [12], and integrated sensing and communication (ISAC) [26].
@@
Nonetheless, the promising performance of FAS relies on the availability of CSI [10, 36, 4, 23], and the high-resolution nature of FAS also leads to substantial overhead in acquiring CSI, including pilot overhead, feedback latency, and energy consumption[13, 8]. In recent years, several studies explored the integration of physical modeling and data-driven approaches for CSI prediction [5, 11, 14]. For instance, [16] proposed an inference algorithm to enable high-accuracy extrapolation in a massive FAS using only a small number of observed ports. Additionally, in [32], an extrapolation scheme based on masked language models, which reconstructs the full channel from a limited number of observed port CSI, was proposed. Going one step further, the approach in [37] integrates the diffusion mechanism of graph neural networks with an asymmetric masked autoencoder architecture, significantly improving both extrapolation accuracy and flexibility in high-resolution FAS settings. Prior to these, there were attempts that exploited channel sparsity [21] and used Bayesian approaches [33]. CSI prediction for multiuser FAS using deep learning was also reported in [27].
However, the existing research shows limitations in reliable CSI acquisition for FAS. Due to the extremely dense deployment of candidate ports, full-port CSI acquisition via sequential switching incurs prohibitive latency and pilot overhead, making extrapolation from a highly limited number of observed ports inevitable. The core challenge lies in reconstructing high-resolution spatial correlation from sparse and noise-contaminated observations, under which conventional learning-based models are prone to noise overfitting and poor generalization to unobserved ports.
Motivated by these challenges, we propose CANet, a deep learning framework specifically designed for robust channel extrapolation in FAS. CANet adopts a synergistic local-to-global architecture with multi-domain regularization, integrating a context-adaptive block (CAB) for reliable local feature fusion, a cross-scale contextual attention (CSCA) mechanism for global feature completion, and a ConvNeXt v2 backbone. In addition, a spatial amplitude perturbation strategy and a Fast Fourier Transform (FFT)-based loss are introduced to enhance robustness against noise by enforcing frequency-domain consistency.
@@ II Channel Extrapolation for FAS
Consider a slow fluid antenna multiple access (FAMA) downlink system[27] where a base station (BS) with $M_{t}$ fixed-position antennas (FPAs) communicates with $U_{t}$ user terminals on the same time-frequency resource unit. Each user terminal employs a 2D FAS with $N_{s}=N_{x}\times N_{y}$ switchable positions (or ports), which are evenly distributed over a planar surface of physical dimensions $W_{s}=W_{x}\times W_{y}$, where each user activates only one port in each coherence block. In a 3D rich scattering environment, the spatial correlation between port $(n_{x}^{s},n_{y}^{s})$ and port $(\tilde{n}_{x}^{s},\tilde{n}_{y}^{s})$ is given by
$\begin{array}[]{l}J_{\left(n_{x},n_{y}\right),\left(\tilde{n}_{x},\tilde{n}_{y}\right)}\\ \qquad=j_{0}\left(\frac{2\pi}{\lambda}\sqrt{\left(\frac{\left|n_{x}-\tilde{n}_{x}\right|}{N_{x}-1}W_{x}\right)^{2}+\left(\frac{\left|n_{y}-\tilde{n}_{y}\right|}{N_{y}-1}W_{y}\right)^{2}}\right),\end{array}$ (1)
where $j_{0}(\cdot)$ denotes the spherical Bessel function of the first kind, which can be approximated by $\mathrm{sinc}(x)={\sin x}/{x}$, and $\lambda$ represents the carrier wavelength. The same spatial correlation matrix $\mathbf{J}_{s}=\mathbf{U}_{s}\mathbf{\Lambda}_{s}\mathbf{U}_{s}^{H}$ is shared by all users, where $\mathbf{U}_{s}\in\mathbb{C}^{N_{s}\times N_{s}}$ and $\mathbf{\Lambda}_{s}\in\mathbb{C}^{N_{s}\times N_{s}}$ denote the eigenvector matrix and the diagonal eigenvalue matrix, respectively.
For a given user, the downlink channel over all FAS ports and all FPAs can be modeled as
$\mathbf{g}=\delta\,\mathbf{U}_{s}\sqrt{\mathbf{\Lambda}_{s}}\mathbf{G},$ (2)
where $\mathbf{g}\in\mathbb{C}^{N_{s}\times M_{t}}$ denotes the small-scale fading channel matrix, $\mathbf{G}\in\mathbb{C}^{N_{s}\times M_{t}}$ consists of independent complex Gaussian random variables with zero mean and variance $1/2$ per real dimension, and $\delta$ characterizes the large-scale path loss.
For the $u$-th user terminal, $u\in\{1,\dots,U_{t}\}$, its channel realization over all FAS ports and BS antennas is denoted by $\mathbf{g}_{u}\in\mathbb{C}^{N_{s}\times M_{t}}$, which follows the same statistical model as in (2). As demonstrated in [1], if the position-to-channel mapping is bijective in a static FAS environment (with fixed geometry, port locations, and scattering conditions), then a channel-to-channel mapping also exists. In other words, once the environment is fixed, each FAS port position almost surely corresponds to a unique downlink channel vector, and vice versa. Unleashing the full potential of FAS necessitates the acquisition of $\mathbf{g}_{u}$. Conventional pilot-based approaches to acquire $\mathbf{g}_{u}$ cause excessive overhead [37]. To this end, we propose channel extrapolation to acquire the CSI of a large set of ports $\widehat{\mathbf{g}}_{u,\mathcal{A}}$ using a small set of observable ports $\mathbf{g}_{u,\mathcal{B}}$, which is expressed as:
$\widehat{\mathbf{g}}_{u,\mathcal{A}}=f(\mathbf{g}_{u,\mathcal{B}}),$ (3)
where f()f(*) is the channel extrapolation function. The aim of the extrapolation is to minimize the expected normalized mean-square-error (NMSE) between the estimated CSI $\widehat{\mathbf{g}}_{u,\mathcal{A}}$ and the ground truth CSI ${\mathbf{g}}_{u,\mathcal{A}}$, which is expressed as:
$\min_{f(*)}\mathbb{E}\left(\frac{|\widehat{\mathbf{g}}_{u,\mathcal{A}}-{\mathbf{g}}_{u,\mathcal{A}}|^{2}}{|{\mathbf{g}}_{u,\mathcal{A}}|^{2}}\right).$ (4)
III Proposed Method
@@ III-A Masking Strategy
We denote the ground truth CSI as $\mathbf{g}\in\mathbb{C}^{N_{s}\times M_{t}}$, and construct a tensor $\mathbf{u}\in\mathbb{R}^{2M_{t}\times N_{y}\times N_{x}}$ to represent the concatenated real and imaginary components of the CSI across a 2D fluid antenna grid of size $N_{x}\times N_{y}$. To simulate partial observations, a subset of antenna ports is randomly masked along the spatial grid, with a masking ratio typically ranging from $80\%$ to $95\%$. The resulting masked tensor is denoted as $\tilde{\mathbf{u}}\in\mathbb{R}^{2M_{t}\times N_{y}\times N_{x}}$, in which the unobserved positions are replaced with a distinct negative value (e.g., 10-10) to make them distinguishable for the network. Additionally, we define a binary indicator matrix $\mathbf{E}_{\text{flag}}\in\{0,1\}^{N_{y}\times N_{x}}$, where a value of $1$ indicates an observed port and 0 is a masked (unobserved) port. No prior knowledge of the channel is used to select ports for extrapolation. Instead, a random masking strategy is employed to ensure flexibility and enhance the generalization capability of the model.
@@ III-B Model Architecture
The architecture of CANet is designed to exploit the spatial continuity and spectral redundancy of FAS. It leverages a local-to-global inductive bias to reconstruct CSI from sparse observations, as detailed below.
@@ III-B1 CAB
To tackle the challenge of irregular sparse observations inherent in FAS, we design the CAB to precisely extract effective information within the receptive field. Standard convolution treats all pixels equally, making it difficult to distinguish valid signals from the values at unobserved positions. To overcome this, CAB employs a context-aware mechanism that dynamically identifies and weighs valid features from observed ports, ensuring accurate local feature extraction.
@@
where $\left\langle\cdot,\cdot\right\rangle$ denotes the composite operation of dot-product with softmax. This self-attention mechanism empowers the model to focus on context information pertinent to the current pixel, thereby enhancing the capability of features to capture and discriminate CSI. In parallel, within the convolutional branch, we employ another set of convolution kernels $w^{c}$ to extract local channel features $x^{c}$ by
$x^{c}=\sum_{i\in\mathcal{B}_{o}}w_{i}^{c}x_{i}.$ (7)
@@ III-B2 CSCA
Considering the finite aperture of FAS, multipath propagation creates specific geometric constraints and long-range correlations across the antenna surface. We introduce CSCA to capture these global dependencies that local convolutions might miss. By facilitating cross-scale contextual attention, it promotes global consistency of the extrapolated channel with the underlying propagation geometry over the finite aperture. We first reduce the input feature channels to half of the original before feeding them into two self-attention modules. One branch directly undergoes self-attention at the same scale, while the other duplicates the feature map and down-samples it to half of the original resolution. The downsampled feature map is then processed along with the original-scale features within the self-attention mechanism. At each self-attention stage, patches are extracted from the input features, and the cosine similarity between patches in the missing region and those in the external region is computed to establish cross-region feature correlations. Formally, let $q_{i}$ denote the $i$-th patch from the masked region $\psi^{l}$, and $k_{j}$ denote the $j$-th patch from the observed region $\phi^{l}$. The cosine similarity is computed as
@@ III-B3 ConvNeXt v2 with dropout regularization
After the initial completion by CAB and CSCA, we employ the ConvNeXt v2 [31] backbone to perform deep feature refinement and learn robust high-level channel structures. ConvNeXt is known for its strong representation capabilities, and ConvNeXt v2 further enhances this by introducing the global response normalization (GRN) module, which enables more expressive feature learning. In our design, we employ the ConvNeXt v2 block integrated with dropout regularization to prevent overfitting and ensure generalization under diverse noise conditions. By randomly dropping a subset of neurons during training, dropout alters the effective network structure in each training epoch, leading to a degree of statistical independence across different training instances, thereby reducing prediction variance and enhancing generalization under diverse noise conditions.
@@ III-B4 Spatial amplitude perturbation
Since FAS channel extrapolation is highly sensitive to measurement noise and multipath variations, we introduce a Spatial Amplitude Perturbation module to enforce robustness by simulating spectral uncertainty. To enhance the model’s robustness, we propose a spatial amplitude perturbation strategy at the input stage, inspired by frequency perturbation techniques in the image domain [18, 35]. Specifically, the input CSI features are first zero-mean normalized to remove the direct current (DC) component, then transformed into the frequency domain via FFT. The amplitude spectrum is then perturbed in a stochastic manner based on a predefined probability, effectively simulating potential frequency interference and noise. This approach improves the model’s adaptability to unseen distributions:
@@
where $\tilde{\mathbf{z}}$ is the spatially zero-mean feature, $\hat{\mathbf{z}}$ is the final perturbed feature after reconstruction, $\bar{\mathbf{z}}$ denotes the spatial mean component removed in (12), $\mathcal{F}(\cdot)$ and $\mathcal{F}^{-1}(\cdot)$ denote the 2D discrete Fourier transform and its inverse. The perturbation mask $\bm{\delta}\in\{0,1\}^{H\times W}$ above is sampled from a Bernoulli distribution with probability $\mu$, and the amplitude coefficients $\bm{\xi}\in[0,1]^{H\times W}$ are sampled from a normal distribution $\mathcal{N}(0,1)$. The hyperparameters $\mu$ and $\gamma$ control the perturbation probability and maximum strength, respectively. Only the amplitude of each frequency component is perturbed, while the phase is fixed to maintain structural consistency.